# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

### v2

@@ Abstract
The adoption of processing-in-memory (PiM) architectures has been gaining momentum because they provide high performance and low energy consumption by alleviating the data movement bottleneck. Yet, the security of such architectures has not been thoroughly explored. The adoption of PiM solutions provides a new way to directly access main memory, which can be potentially exploited by malicious user applications. We show that this new way to access main memory opens opportunities for high-throughput timing attack vectors that are hard-to-mitigate without significant performance overhead.
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of PiM architectures to establish covert and side channels. IMPACT enables high-throughput communication and private information leakage. To achieve this, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric main memory and cache-based timing attacks and (ii) leverages the intrinsic parallelism of PiM operations. First, we showcase two covert-channel attack variants that run on the host CPU and leverage PiM architectures to gain direct and fast access to main memory and establish high-throughput communication covert channels. Second, we showcase a side-channel attack on a DNA sequence analysis application that leaks the private characteristics of a user’s sample genome by leveraging PiM operations. Our results demonstrate that (i) our covert channels achieve up to 14.16 Mb/s communication throughput, which is 5.41$\times$ faster than the state-of-the-art main memory-based covert channels, and (ii) our side-channel attack allows the attacker to determine the properties of a sample genome at a throughput of 7.5 Mb/s with 96% accuracy. We discuss and evaluate several countermeasures for IMPACT to enable secure and robust PiM architectures.
@@ 1 Introduction
Data movement between the main memory and the processor is a prevalent bottleneck present in today’s computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Hence, both industry and academia have invested significant efforts to bring computation closer to where data resides, creating a paradigm called processing-in-memory (PiM) [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. PiM solutions provide significant performance improvement and energy efficiency by alleviating data movement.
PiM solutions generally take one of two approaches: (i) processing-near-memory (PnM), and (ii) processing-using-memory (PuM). First, PnM techniques offload computation to compute units near memory arrays (e.g., in the logic layer of 3D-stacked memory or within the memory chip) [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]. During recent years, multiple real-world PnM systems have emerged, such as UPMEM [43], AxDIMM [41], AiM [44], HB-PNM [45] and FIMDRAM [46]. At the same time, several academic works propose using general-purpose programmable PiM devices where the user application offloads instructions to the compute units inside the memory at fine-granularity [47] (e.g., PiM-enabled instructions [48]) or coarse granularity (i.e., full program [20, 49]). Second, PuM techniques propose exploiting the analog properties of memory operations and performing specialized PiM operations. Prior works improve the performance of workloads by providing bulk data copy and initialization, bitwise operations, and arithmetic operations using memory [50, 51, 52, 53, 54, 55, 56, 49, 57, 58, 59, 60, 61, 62, 63, 64]. However, to the best of our knowledge, no prior work evaluates the security of PiM architectures against timing attacks.
@@
Analyzing and evaluating the security of PiM architectures is critical to prevent widespread vulnerabilities as PiM architectures become widely available. After a thorough analysis of existing PiM architectures, in this work, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons: (i) PiM provides direct access to main memory , and (ii) a large set of existing practical defense mechanisms are not applicable in PiM systems.
1. Direct Access to Main Memory. Timing attacks in PiM architectures become easier to exploit with the direct memory access provided to user space applications by these architectures. The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the user space applications to bypass caches more easily (i.e., with fewer memory accesses), and (ii) PuM architectures provide the ability to offload specialized PiM instructions (e.g., bulk data copy and initialization, and bitwise/arithmetic operations) to the main memory that completely bypass the deep cache hierarchies of modern processors.
2. Hard-to-Mitigate with Practical Defenses. A fundamental solution to memory-based timing attacks is to eliminate the timing channel itself. Prior mitigations explore fundamental solutions such as memory partitioning and enforcing constant-time memory accesses. However, these mitigations incur significantly high performance overheads (as shown in Section 6). Therefore, practical mitigations and detection mechanisms restricting access to timing channels are explored. Several prior works [65, 66, 67, 68] propose cache management-based methods such as restricting the access to cache management instructions (e.g., clflush instruction in x86 systems [65]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [66, 67, 68]. These mechanisms are inapplicable to PiM-based attacks because these attacks completely bypass the cache hierarchy.
Our goal is to develop a set of high-throughput and hard-to-mitigate timing side and covert-channel attacks in PiM architectures. We expect and hope that IMPACT will bring attention to security issues in emerging PiM architectures and result in security awareness in PiM architecture designs.
In this work, we introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverages PiM architectures to gain direct and fast access to main memory, establish high-throughput communication covert channels and leak secrets of concurrently running applications. IMPACT leverages the direct access, provided by PiM operations, to the shared DRAM row buffer and achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing cover-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations.
@@
Key Challenge and Limitations of Prior Work. For successful main memory-based timing side- and covert-channel attacks, attacker processes (or sender and receiver processes for covert-channel attacks) need to access the main memory with high frequency to observe or change the states of the shared main memory structures. In traditional architectures, it is difficult to access the main memory with high frequency due to the deep cache hierarchies that filter the memory requests of attacker processes.
We analyze (Section 3.2) the effectiveness of the attack primitives that bypass caches and provide direct access to the main memory to leverage the main memory timing channel: clflush instruction, eviction sets, (remote) direct memory access, (R)DMA, requests, non-temporal memory hints (e.g., movnti instruction [69]), and PiM operations. None of these primitives used in traditional architectures provide fast and reliable direct memory access at the same time. In contrast, PiM operations provide fast and reliable direct memory access to the user space applications. Therefore, PiM architectures can be more amenable to high-throughput timing side- and covert-channel attacks compared to traditional architectures.
@@
IMPACT Covert-Channel Attack Variants. We showcase two different timing covert-channel attack variants by exploiting PnM and PuM techniques. First, we build a PnM-based covert channel, IMPACT-PnM, that uses PiM-enabled instructions [48] to directly access the main memory. The sender bypasses the locality monitor employed by the system and offloads PiM-enabled instructions for computation directly inside the main memory. The receiver measures the time it takes to access a specific set of rows to retrieve the message encoded with row buffer hits and conflicts. IMPACT-PnM achieves 12.87 Mb/s communication throughput, which is up to 5.79$\times$ faster than the state-of-the-art main memory-based covert-channel attack [70]. Second, we present IMPACT-PuM, a PuM-based covert channel that leverages the inherent parallelism in main memory using bulk copy and initialization techniques, specifically RowClone [55], to quickly transmit messages through the shared row buffer states. This way, IMPACT-PuM transmits 16-bit large messages 14$\times$ faster than IMPACT-PnM and achieves communication throughput of 14.16 Mb/s (up to 6.38$\times$ faster than the state-of-the-art [70]).
@@
IMPACT Side-Channel Attack. We showcase a side-channel attack on a DNA sequence analysis application: DNA read mapping. Preserving the privacy of DNA sequences is crucial, especially when the sequences are used for medical purposes. DNA read mapping finds the matching locations of short DNA fragments of the sample genome in the reference genome, which acts as a template genome to construct the complete genome of a given sample. Our side-channel attack leverages PiM operations and leaks the private characteristics of a user’s sample genome by observing the memory access patterns of the DNA read mapping application. Our evaluation demonstrates that our side-channel attack enables determining the properties of a sample genome at a throughput of 7.5 Mb/s with 96% accuracy.
@@
Defense Mechanisms. Based on our security analysis, we discuss three and evaluate two potential defense mechanisms that can mitigate IMPACT attacks by eliminating the timing channel. We discuss the performance versus security trade-offs of these countermeasures. We demonstrate that the two best-performing defense mechanisms lead to high performance overheads of 25% and 16% on average across five data-intensive workloads.
@@
  • We present IMPACT, the first timing attack in PiM systems. We leverage the direct memory access enabled by PiM operations to establish high-throughput and reliable timing covert channels.
  • We showcase two IMPACT variants by exploiting PnM and PuM techniques: IMPACT-PnM and IMPACT-PuM. IMPACT-PnM allows the sender and the receiver to exchange messages by using PiM instructions. IMPACT-PuM further leverages the parallelism provided by bulk copy PuM techniques to improve the throughput. We extensively evaluate both IMPACT-PnM and IMPACT-PuM and show that they significantly improve communication throughput compared to the state-of-the-art main-memory-based covert channel.
  • We showcase an IMPACT side-channel attack on DNA read mapping by leveraging PnM instructions. We show that our side-channel attack enables the attacker to leak sensitive information on the victim’s sample genome with high throughput and a low error rate.
  • We evaluate two defense mechanisms to protect the system against IMPACT attacks caused by the adoption of PiM solutions. We discuss the trade-offs of these mechanisms to aid the adoption of robust and secure PiM architectures.
@@ 2.1 DRAM Organization and Operation
Organization. Figure 1 illustrates a DRAM module’s hierarchical organization. A DRAM chip is organized into multiple DRAM banks, each of which shares a common I/O circuitry in a DRAM bank [71, 72, 73, 74, 75]. DRAM cells are organized as a two-dimensional array to form a bank and are addressed by its row and column [76, 77, 22, 78]. The DRAM cell is connected to a bitline that is used to access the data stored in the cell and a wordline that controls access to the cell. Each DRAM cell in a DRAM row is connected to a common wordline via its access transistors. A bitline connects a column of DRAM cells to a DRAM sense amplifier to read or write data. A row of sense amplifiers is called a row buffer [56, 79, 80, 81, 82, 55]. At the lowest level of the hierarchy, a single DRAM cell comprises 1) a storage capacitor that stores a single bit of data encoded using the charge level in the capacitor and 2) an access transistor that is used to read from and write to the storage capacitor [83, 84, 85, 86, 87].
@@
Operation. The memory controller serves memory access requests. To serve memory requests, the memory controller issues DRAM commands, e.g., row activation ($ACT$), bank precharge ($PRE$), data read ($RD$), data write ($WR$), and refresh ($REF$). Between consecutive DRAM commands, the memory controller must obey standardized timings to ensure correct DRAM operation. For read and write requests, the memory controller first issues an $ACT$ command to the corresponding bank, which opens the corresponding row. The data in the row is fetched to the row buffer, where it is temporarily stored. The data in the row buffer becomes accessible by a $RD$/$WR$ command when the bitline and cell voltage are restored to a certain level. Hence, the memory controller should wait at least the activation latency ($t_{RCD}$) between an $ACT$ and a $RD$/$WR$ command. All changes are propagated to the DRAM cells in the open row. Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [88]. However, in the case where the target row is not in the row buffer, the memory controller issues a $PRE$ command to close the already open row and issues an $ACT$ command to fetch the target row, creating a row conflict/miss. The time interval for issuing 1) a $PRE$ after an $ACT$ must be at least restoration latency ($t_{RAS}$) and 2) an $ACT$ after a $PRE$ must be at least the precharge latency ($t_{RP}$).
@@ 2.2 Processing-in-Memory
Processing-in-memory (PiM) is a promising paradigm that moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. There are two main approaches to PiM: 1) processing-near-memory (PnM) and 2) processing-using-memory (PuM).
Processing Near Memory (PnM). In processing-near-memory (PnM) computation is placed near the memory arrays. PnM takes advantage of the computation capability of conventional memory controllers or the logic layer(s) of the 3D-stacked memory technologies [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]. PnM logic can execute portions of applications (e.g., instructions [48, 32], functions [28], whole application threads [20, 49]) depending on the PnM architecture. PnM computation units typically employ small caches due to 1) area constraints in the PnM logic and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications. Therefore, various PnM systems [32, 48, 28] execute the memory-intensive regions of applications on the PnM computation units, while executing the more cache-friendly regions of applications on the host CPU with a deeper cache hierarchy.
Processing Using Memory (PuM). Processing Using Memory (PuM) expands the functionality of the memory arrays with minimal changes so that they can perform computation [50, 51, 52, 53, 54, 55, 56, 49, 57, 58, 59, 60, 61, 62, 63, 64]. This approach leverages the properties and operational principles of memory cells and arrays to perform computation, without the need for dedicated logic unit added to the memory system. PuM architectures take advantage of the high bandwidth of the memory arrays.
2.3 Microarchitectural Timing Attacks
Microarchitectural timing attacks exploit the observable state changes in microarchitecture to build covert channel attacks that can propagate messages between processes and leak secrets of concurrently running processes. These attacks measure the time to complete an operation that depends on the shared structures in the system (e.g., accessing a memory location by going through the shared caches) to infer the states of these shared structures.
Main Memory Timing Attacks. Existing main memory timing attacks exploit several shared structures in the main memory to leak data, and one such structure is the row buffer. The row buffer acts as an internal cache inside DRAM banks. Thus, an attacker can understand if a row is recently accessed (i.e., already in the row buffer) or not by measuring access to that row. Based on this observation, prior works propose several attacks, revealing the DRAM address mapping and leaking keystrokes. DRAMA [70] (i) builds a covert channel that exchanges messages between processes using the row buffer conflicts and hits, and (ii) showcases a side-channel attack that can leak keystrokes in a real system. Similarly, many works [70, 89, 90] exploit row buffer-based timing differences to determine DRAM address mappings. Other attacks based on the main memory exploit either the memory bus [91], or observations of latencies incurred by other memory-related operations such as memory deduplication [92, 93], and memory compression [94, 95, 96, 97, 98, 99, 100, 101] and decompression [102].

### v3

@@ Abstract
The adoption of processing-in-memory (PiM) architectures has been gaining momentum because they provide high performance and low energy consumption by alleviating the data movement bottleneck. Yet, the security of such architectures has not been thoroughly explored. The adoption of PiM solutions provides a new way to directly access main memory, which malicious user applications can potentially exploit. We show that this new way to access main memory opens opportunities for high-throughput timing attack vectors that are hard-to-mitigate without significant performance overhead.
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of PiM architectures to establish covert and side channels. IMPACT enables high-throughput communication and private information leakage by exploiting the shared DRAM row buffer. To achieve high throughput, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric main memory and cache-based timing covert and side-channel attacks and (ii) leverages the intrinsic parallelism of PiM operations. We showcase two applications of IMPACT. First, we build two covert-channel attack variants that run on the host CPU and leverage different PiM approaches (i.e., processing-near-memory and processing-using-memory) to gain direct and fast access to main memory and establish high-throughput communication covert channels. Second, we showcase a side-channel attack that leaks private information of concurrently running victim applications that are accelerated with PiM. Our results demonstrate that (i) our covert channels achieve 12.87 Mb/s and 14.16 Mb/s communication throughput, respectively, which is up to 4.91$\times$ and 5.41$\times$ faster than the state-of-the-art main memory-based covert channels, and (ii) our side-channel attack allows the attacker to leak secrets with a low error rate. To avoid such covert and side channels in emerging PiM systems, we propose and evaluate three defenses that fundamentally eliminate the timing channel.
@@ 1 Introduction
Data movement between the main memory and the processor is a critical bottleneck present in today’s computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Hence, both industry and academia have invested significant efforts to bring computation closer to where data resides, creating a paradigm called processing-in-memory (PiM)  [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. PiM solutions provide significant performance improvement and energy efficiency by reducing data movement.
PiM solutions take either of two approaches: (i) processing-near-memory (PnM), where the computation happens in a computation unit placed on the same chip as the memory array [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42], which also appears in real-world products (e.g., UPMEM [43], AxDIMM [41], AiM [44], HB-PNM [45] and FIMDRAM [46]) and (ii) processing-using-memory (PuM) where the memory array performs the computation by leveraging the analog computation capability of DRAM cells [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. Prior works on PnM and PuM clearly demonstrate that these emerging architectures promise great performance and energy efficiency benefits. In addition, the existence of real products that implement such techniques [43, 41, 44, 45, 46] indicates that these techniques are likely to be used in many real systems in the near future. Therefore, it is important to investigate the security aspect of such techniques and implement countermeasures to avoid potential widespread security vulnerabilities that their deployment can cause. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
@@
In this work, we analyze existing PiM architectures and show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based attacks that are hard-to-mitigate without significant performance overheads. This is possible due to two main reasons: (i) PiM architectures provide direct access to main memory, which is essential for their performance and energy efficiency benefits, and (ii) a large set of existing practical defense mechanisms are not applicable in these PiM architectures.
@@
1. Direct Access to Main Memory. High-throughput main memory-based timing side- and covert-channel attacks need to access main memory with high frequency to observe or change the states of shared main memory structures (e.g., DRAM row buffers). In today’s systems, it is difficult to access main memory with high frequency due to the deep cache hierarchies that filter the memory requests of attacker processes. [^1] We provide analysis as to why prior attack primitives that bypass caches do not provide fast and reliable direct main memory access in §3.2.
The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the user space applications to bypass caches more easily (i.e., with fewer memory accesses), and (ii) PuM architectures provide the ability to offload specialized PiM instructions (e.g., bulk data copy and initialization, and bitwise/arithmetic operations) to the main memory that completely bypass the deep cache hierarchies of modern processors. As a result, PiM architectures provide user space applications with fast and reliable direct main memory access that can be exploited by attackers for high-throughput timing covert- and side-channel attacks.
2. Hard-to-Mitigate with Practical Defenses. A fundamental solution to memory-based timing attacks is to eliminate the timing channel itself. Prior defenses explore fundamental solutions such as memory partitioning and enforcing constant-time memory accesses. However, these defenses incur significantly high performance overheads (as shown in Section 6). Therefore, several prior works [63, 64, 65, 66] explore practical defenses and detection mechanisms that restrict access to timing channels with cache management-based methods, such as restricting the access to cache management instructions (e.g., clflush instruction in x86 systems [63]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [64, 65, 66]. These mechanisms are inapplicable to PiM-based attacks because these attacks completely bypass the cache hierarchy.
Based on our analysis, we introduce IMPACT, a set of high-throughput In-Memory Processing-based timing Attacks that leverage the direct and fast main memory accesses enabled by PiM architectures. Using the direct access via PiM operations, IMPACT observes and exploits the shared DRAM row buffer. We demonstrate that IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations. We showcase this finding on two covert- and one side-channel attacks. To protect future systems against IMPACT covert and side channels, we propose three defense mechanisms that fundamentally eliminate the channel itself.
@@
IMPACT-PnM Covert Channel Variant. We build a PnM-based covert channel that uses PiM-enabled instructions [67] to directly access the main memory. The sender bypasses the locality monitor employed by the system and offloads PiM-enabled instructions for computation directly inside the main memory. The receiver measures the time it takes to access a specific set of rows to retrieve the message encoded with row buffer hits and conflicts. IMPACT-PnM achieves 12.87 Mb/s communication throughput, which is up to 4.91$\times$ faster than the state-of-the-art main memory-based attack [68].
IMPACT-PuM Covert Channel Variant. We build a PuM-based covert channel that leverages the inherent parallelism in main memory using bulk copy and initialization techniques, specifically RowClone [52], to quickly transmit messages through the shared row buffer states. This way, IMPACT-PuM transmits 16-bit large messages 14$\times$ faster than IMPACT-PnM and achieves communication throughput of 14.16 Mb/s (up to 5.41$\times$ faster than the state-of-the-art [68]).
@@
IMPACT Side-Channel Attack. We build a PnM-based IMPACT side-channel attack that exploits PiM operations to leak private information belonging to a concurrently running application based on its memory access characteristics. We showcase this side-channel attack on read mapping, one of the fundamental applications in DNA sequence analysis where preserving the privacy of the human genomic data is crucial [69, 70, 71, 72, 73]. Our side-channel attack leverages PiM operations and leaks the private characteristics of a user’s sample genome by observing the memory access patterns of the read mapping application. Our evaluation demonstrates that our side-channel attack leaks the properties of a sample genome at a throughput of 7.57 Mb/s with 96% accuracy.
@@
Defense Mechanisms. Based on our security analysis, we discuss and evaluate potential defense mechanisms that can mitigate IMPACT attacks by eliminating the timing channel. We discuss the performance versus security trade-offs of these defenses. We demonstrate that the two best-performing defense mechanisms lead to high performance overheads of 26% and 15% on average across five data-intensive workloads.
@@
  • To the best of our knowledge, this is the first work that analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks. We introduce IMPACT, a new set of high-throughput main memory-based timing attacks that leverage PiM operations to gain fast and reliable access to main memory.
  • We showcase two IMPACT covert channel variants, IMPACT-PnM and IMPACT-PuM that exploit PnM and PuM techniques, respectively. First, IMPACT-PnM allows the sender and the receiver to exchange messages using PiM-enabled instructions. Second, IMPACT-PuM further leverages the parallelism provided by bulk copy PuM techniques to improve the communication throughput. We show that both attacks outperform the state-of-the-art main-memory-based covert-channel attacks in terms of communication throughput.
  • We showcase an IMPACT side-channel attack on read mapping by leveraging PnM instructions. We show that our side-channel attack enables the attacker to leak sensitive information on the victim’s sample genome with high throughput and a low error rate.
  • We evaluate and discuss the trade-offs of three defense mechanisms to protect PiM-enabled systems against IMPACT’s covert- and side-channel attacks.
@@ 2.1 DRAM Organization and Operation
Organization. Fig. 1 shows the hierarchical organization of a modern DRAM-based main memory. The memory controller connects to a DRAM module over a memory channel. A module contains one or more DRAM ranks that time-share the memory channel. A rank consists of multiple DRAM chips that operate in lock-step. Each DRAM chip contains multiple DRAM banks that can be accessed independently. A DRAM bank is organized as a two-dimensional array of DRAM cells, where a row of cells is called a DRAM row. A DRAM cell consists of 1) a storage capacitor, which stores one bit of information in the form of electrical charge, and 2) an access transistor, which connects the capacitor to the row buffer through a bitline controlled by a wordline.
@@
Operation. To perform a read or write operation, the memory controller first needs to open a row, i.e., copy the data of the cells in the row to the row buffer. To open a row, the memory controller issues an activation command to a bank by specifying the row address to open. After activation is complete, the memory controller issues either read and write commands to read or write a DRAM word within the activated row. Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [74]. To access data from another DRAM row in the same bank, the memory controller must first close the currently open row by issuing a bank precharge command, creating a row conflict/miss.
@@ 2.2 Processing-in-Memory
Processing-in-memory (PiM) moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. There are two main PiM approaches: 1) processing-near-memory (PnM) and 2) processing-using-memory (PuM). PnM place compute logic near memory arrays (e.g., processing capability in the memory controller and logic layer(s) of 3D-stacked memory [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]). PnM logic can execute application regions (e.g., instructions [67, 32], functions [28], application threads [20, 54]) depending on the design. PnM computation units typically employ small caches due to 1) area constraints and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications. Therefore, various systems [32, 67, 28] offload the memory-intensive application regions on the PnM units, while executing more cache-friendly application regions on the host CPU with a deeper cache hierarchy. PuM uses analog operational properties of the memory circuitry (e.g., DRAM) to enable massively parallel in-memory computation [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
2.3 Main Memory Timing Attacks
Microarchitectural timing attacks exploit the observable microarchitectural state changes to propagate messages between processes and leak secrets. These attacks measure the time to complete an operation that depends on a shared system structure (e.g., accessing a memory location by going through the shared caches) to infer the state of the shared structure.
Main memory timing attacks exploit several shared structures in the main memory. One such structure is the DRAM row buffer that acts as an internal cache inside DRAM banks. An attacker can understand if a row is recently accessed (i.e., already in the row buffer) or not by measuring the time it takes to access that row. Using this observation, DRAMA [68] builds a side-channel attack to leak keystrokes in a real system. Several prior works [68, 75, 76, 77, 78] leak DRAM address mapping functions. Other main memory-based attacks exploit the memory bus [79], and memory deduplication [80, 81] and (de)compression [82, 83, 84, 85, 86, 87, 88, 89, 90] latencies.